Theory and Applications of Categories, Vol. 3, No. 8, 1997, pp. 182–212.

MONADS AND INTERPOLADS IN BICATEGORIES

JÜRGEN KOSLOWSKI

Transmitted by R. J. Wood

ABSTRACT. Given a bicategory, Y , with stable local coequalizers, we construct a bicategory of monads Y -mnd by using lax functors from the generic 0-cell, 1-cell and 2-cell, respectively, into Y . Any lax functor into Y factors through Y -mnd and the 1-cells turn out to be the familiar bimodules. The locally ordered bicategory rel and its bicategory of monads both fail to be Cauchy-complete, but have a well-known Cauchycompletion in common. This prompts us to formulate a concept of Cauchy-completeness for bicategories that are not locally ordered and suggests a weakening of the notion of monad. For this purpose, we develop a calculus of general modules between unstructured endo-1-cells. These behave well with respect to composition, but in general fail to have identities. To overcome this problem, we do not need to impose the full structure of a monad on endo-1-cells. We show that associative coequalizing multiplications suffice and call the resulting structures interpolads. Together with structure-preserving i-modules these form a bicategory Y -int that is indeed Cauchy-complete, in our sense, and contains the bicategory of monads as a not necessarily full sub-bicategory. Interpolads over rel are idempotent relations, over the suspension of set they correspond to interpolative semi-groups, and over spn they lead to a notion of “category without identities” also known as “taxonomy”. If Y locally has equalizers, then modules in general, and the bicategories Y -mnd and Y -int in particular, inherit the property of being closed with respect to 1-cell composition.

Introduction

Part of the original motivation for this work was to better understand, why the bicategory idl of pre-ordered sets, order-ideals, and inclusions inherits good properties from the bicategory rel of sets, relations and inclusions. Of particular interest was closedness with respect to 1-cell composition, also known as the existence of all right liftings and right extensions.

The key observation is that pre-ordered sets can be viewed as monads in rel . Benabou [1] explicitly designed what is now known as lax functors to subsume the notion of monad, in this case as a lax functor from the terminal bicategory 1 to rel . To keep the paper reasonably self-contained, in Section 1 we recall the relevant definitions and establish our notation.

Order-ideals, i.e., relations compatible with the orders on domain and codomain, do not correspond to the kind of morphisms that are usually considered between monads. Since ordinary morphisms in a category correspond to functors with domain 2 , the two-element chain or generic 1-cell, we introduce 1-cells between monads in terms of

Received by the editors 1997 January 22 and, in revised form, 1997 September 24. Published on 1997 October 14 1991 Mathematics Subject Classification : 18D05. Key words and phrases: bicategory, closed bicategory, Cauchy-complete bicategory, lax functor,

monad, module, interpolation property, taxonomy.

\begin{matrix}c \\ ◯ \end{matrix} Jürgen Koslowski 1997. Permission to copy for private use granted.

lax functors with 2 as domain and call them “m-modules”. These later turn out to be bimodules between monads as described in Section 4.1 of Carboni, Kasangian and Walters [3] and also known to Street [9]. Since the inclusions in idl correspond to lax functors from the generic 2-cell, we may interpret idl as “the bicategory of monads over rel ”.

Rosebrugh and Wood [8] mention a basic defect of rel : idempotents need not split. This also is true for idl . Hence they work with the Karoubian envelope, or Cauchy-completion, kar of rel instead. It also has all right liftings and right extensions and happens to be the Cauchy completion of (and hence to contain) idl as well.

How can this situation be generalized to other well-behaved bicategories besides rel ? There are (at least) two aspects to this question. The objects of kar are sets equipped with an idempotent relation. This need not be reflexive but has to be transitive and to satisfy the so-called “interpolation property”, which is just the converse of the inclusion required for transitivity, cf. Example 3.07.(2). Hence one may think of it as a “monad without unit”, but with an extra condition on the multiplication that in the general case has yet to be determined. We chose the name “interpolad” for such a structure. Of course, morphisms between interpolads ought to be suitably weakened versions of m-modules or bimodules. In fact, no extra structure on endo-1-cells is necessary to get a theory of general modules off the ground. In Section 2 we describe this for a fixed bicategory Y with local stable coequalizers. General modules already admit the type of composition, by means of coequalizers, described for bimodules by Carboni et. al [3], and known in various special cases. Unfortunately, endo-1-cells and general modules do not form a bicategory for lack of identity modules.

This prompts us, in Section 3, to look for the minimal structure that needs to be imposed on endo-1-cells and to be preserved by suitable modules in order to guarantee the existence of identities. This leads to the official introduction of interpolads: endo-1-cells with an associative coequalizing multiplication. Together with structure-preserving “imodules” and the obvious 2-cells they form a bicategory Y -int . The extent to which the notions of monad and m-module have been weakened can be made precise: the absolute coequalizers implicit in those notions are replaced by coequalizers, the preservation of which in certain cases we require explicitly. Hence we recover Y -mnd as a not necessarily full sub-bicategory of Y -int .

Section 4 addresses the second aspect of our quest to generalize the relationship between rel , idl and kar : we introduce the notion of Cauchy-completeness for arbitrary bicategories. If the hom-categories are not just ordered, associativity becomes an issue when reasoning about idempotency of 1-cells. Y -int turns out to be Cauchycomplete in this sense, and hence to fully contain the Cauchy-completion of Y . The latter will usually be too small to contain Y -mnd . Although for distributive Y all interpolads arise when splitting idempotent m-modules in Y -int , the Cauchy-completion of Y -mnd usually cannot be realized as a subcategory of Y -int , although there exists an obvious “forgetful” functor.

In Section 5 we return to our original question and consider closedness with respect to 1-cell composition in terms of right extensions and right liftings. Provided that Y has hom-categories with equalizers, we show that the existence of all right extensions in Y implies this property for general modules (no identities are needed for this), as well as for the bicategories Y -mnd and Y -int . In fact, right extensions in Y -mnd are formed just as for general modules (this recovers parts of a result by Street [9]), while in Y -int their construction requires an additional step, the pre- and post-composition with the appropriate identity i-modules. Right liftings can be handled dually.

It is known that the bicategory of monads over spn , the bicategory of sets, spans or matrices of sets, and span morphisms or matrices of functions, is essentially the bicategory of small categories and profunctors, cf. Example 3.07.(4). The bicategory of interpolads over spn seems to be the largest reasonable extension to a bicategory with objects that resemble categories but may fail to have identities. We propose to call them “taxonomies”, even though this term has been introduced by Paré and Wood for a weaker concept (no further conditions on the associative multiplication).

The results of this paper seem to suggest that the unit condition for a binary multiplication may be “less fundamental” than the associativity condition. Indeed, both basic diagrammatic representations of bicategorical data, pasting diagrams and string diagrams, can account for the unit isomorphisms more directly than for the associativity isomorphisms. But this needs to be studied more closely in the future.

Early versions of this paper have been presented at CT95 in Halifax, and at the Isle of Thorns Category Theory meeting, 1996.

1. Bicategories and lax functors

Bicategories and morphisms between them were introduced by Benabou [1]. Recently, Borceux [2] has covered some of this material in a textbook. Roughly speaking, the notion of bicategory arises by replacing the hom-sets of ordinary categories by hom-categories, the composition functions by functors, and by then relaxing the categorical axioms for morphisms to hold “up to coherent isomorphism”. While the idea is straightforward, the actual formulation of “coherent isomorphism” requires some effort. Hence we briefly recall the definitions of bicategory and of lax functor and introduce our preferred notation (cf. Conventions 1.01. and 1.05. below).

1.00. Definition. A bicategory X consists of

• a class \underline{X} of objects or 0-cells

• for all 0-cells A, B a small category ⟨A, B⟩X ; its objects (denoted by single arrows) are called 1-cells, and its morphisms (denoted by double arrows) are called 2-cells; we refer to the composition of 2-cells in ⟨A, B⟩X as serial or vertical composition, and write f [figure: σ ; τ] h for the composite of f [figure: σ] g and g [figure: τ] h; ^{1}

• for all 0-cells A, B, and C a functor ⟨A, B⟩X × ⟨B, C⟩X [figure: ⟨A,B,C⟩c] ⟨A, C⟩X , referred to as parallel or horizontal composition; we concatenate the arguments, e.g., φψ instead of ⟨φ, ψ⟩ (⟨A, B, C⟩c) ; notationally, this binds stronger than “ ; ”; • for all composable 1-cells A [figure: f] B [figure: g] C [figure: h] D a natural associativity isomor-

1 Readers unhappy with the order of composition can obtain a version of this paper in backwards notation from http://www.iti.cs.tu-bs.de/TI-INFO/koslowj/koslowski.html or from the author.

phism f(gh) [figure: ⟨f,g,h⟩𝔞] (fg)h;

• for each 0-cell A a distinguished identity 1-cell in ⟨A, A⟩X , usually identified with the object A, or, if necessary, notationally disambiguated as 1_{A};

• for each 1-cell A [figure: f] B natural unit isomorphisms Af [figure: f𝔲_{⋄}] f and fB [figure: f𝔲^{⋄}] f

subject to the following coherence axioms:

(B0) for all composable 1-cells A [figure: f] B [figure: g] C [figure: h] D [figure: k] E we have

[figure: f((gh)k); f(g(hk)); ^{(fg)(hk)}; ((fg)h)k; (f(gh))k; f((⟨g,h,k⟩𝔞)); ⟨f,gh,k⟩𝔞; ((⟨f,g,h⟩𝔞))k; ⟨f,g,hk⟩𝔞; ⟨fg,h,k⟩𝔞]

For later reference we call this property of the composition functors and the associativity isomorphisms the essential associativity of horizontal composition.

(B1) for all composable 1-cells A [figure: f] B [figure: g] C we have

[figure: (fB)g; f(Bg); fg; ⟨f,B,g⟩𝔞; (f𝔲^{⋄})g; f(g𝔲_{⋄})]

If there is no danger of confusion, we may drop the arguments to 𝔞, 𝔲_{⋄}, and 𝔲^{⋄}. If all these structural 2-cells are identities, we call X a 2-category.

dual notions: In the bicategories X^{op} and X^{co} , the 1-cells and the 2-cells are reversed, respectively. I.e., ⟨A, B⟩X^{op} = ⟨B, A⟩X , and ⟨A, B⟩X^{co} = (⟨A, B⟩X )^{op} .

1.01. Convention. We usually denote 0-cells by upper case letters, 1-cells with lower case letters, and 2-cells with lower case Greek letters. Exceptions are the structural associativity and unit isomorphisms, which we distinguish from “ordinary” 2-cells by means of bold Fraktur letters. The identity 2-cell of A [figure: r] B is usually identified with r, or, if necessary, notationally disambiguated as id_{r}. Diamonds ⋄ as subscripts and superscripts always indicate left and right actions, respectively.

1.02. Remark. The coherence axioms guarantee, among other things, that

(0) for every object A the unit isomorphisms AA [figure: A𝔲_{⋄}] A and AA [figure: A𝔲^{⋄}] A agree; this justifies dropping the diamond, i.e., we define A𝔲 := A𝔲_{⋄} = A𝔲^{⋄}; (1) for any two composable 1-cells A [figure: f] B [figure: g] C we have

[figure: (Af)g; A(fg); fg; ⟨A,f,g⟩𝔞; (f𝔲_{⋄})g; _{(fg)𝔲_{⋄}}; and; (fg)C; f(gC); fg; ⟨f,g,C⟩𝔞; _{(fg)𝔲^{⋄}}; f(g𝔲^{⋄})]  (1-00)

1.03. Examples. (0) Every “ordinary” category can be viewed as a bicategory with discrete hom-categories. Since it will always be clear from the context, which interpretation is intended, we will use the same name, e.g., set , in both cases.

(1) One of the simplest bicategories is rel with sets as objects, binary relations as 1-cells and inclusions as 2-cells. 1-cell composition is the usual relation product. Since the hom-categories are partially ordered, this is in fact a 2-category.

(2) A very important bicategory, bim , has unitary rings as objects. A 1-cell from A to B is an ⟨A, B⟩-bimodule, i.e., an abelian group R equipped with the structures of a left-A-module and a right-B -module subject to the compatibility condition

a(rb) = (ar)b

for all a ∈ A, r ∈ R and b ∈ B, cf. Examples 3.07.(0) and (1). 2-cells between ⟨A, B⟩-bimodules are group homomorphisms that are A-linear on the left and B - linear on the right. Bimodules are composed by means of their tensor product. For a commutative unitary ring R, the hom-category ⟨R, R⟩bim contains the familiar category of R-modules as a full subcategory. Hence bim is a true bicategory that is not a 2-category. Much of the terminology derives from this particular example.

(3) For any object A of a bicategory X the hom-category ⟨A, A⟩X carries the structure of a monoidal category, the tensor product being the composition of 1-cells. In fact, every monoidal category M may be viewed as the hom-category of a bicategory with one object {∗}. The Australian school has coined the term suspension of M for the resulting bicategory M Σ that satisfies ⟨∗, ∗⟩ (M Σ) = M .

1.04. Definition. For bicategories X and Y , a lax functor X [figure: F] Y consists of

• an object-function or carrier \underline{X} [figure: F] \underline{Y} ;

• a family of functors ⟨A, B⟩X [figure: ⟨A,B⟩F] ⟨AF, BF ⟩Y , where A and B are X - objects; we usually write fF instead of f⟨A, B⟩F ;

• a family of natural 2-cells (fF ) (gF ) [figure: ⟨f,g⟩𝔮] (fg)F , where A [figure: f] B [figure: g] C are composable 1-cells in X ;

• a family of natural 2-cells AF [figure: A𝔡] AF , where A is an X -object

subject to the following coherence axioms:

(L0) for all composable 1-cells A [figure: f] B [figure: g] C [figure: h] D in X we have

[figure: (fF ) ((gF ) (hF )); ^{(fF )}^{(⟨g,h⟩𝔮)}; ⟨fF,gF,hF⟩ 𝔞; (fF ) ((gh)F ); ⟨f,gh⟩𝔮; (f(gh))F; (⟨f,g,h⟩𝔞)F; (_{(fF )}_{(gF )})_{(hF )}; _{(⟨f,g⟩𝔮)}_{(hF )}; _{((fg)F )}_{(hF )}; ⟨fg,h⟩𝔮; ((fg)h)F]

(L1) for each 1-cell A [figure: f] B in X we have

[figure: (AF) (fF ); ^{(fF )}𝔲_{⋄}; _{(A𝔡)}_{(fF )}; fF; _{(AF )}_{(fF )}; ⟨A,f⟩𝔮; (Af)F; (f𝔲_{⋄})F; and; (fF ) (BF); ^{(fF )}𝔲^{⋄}; _{(fF )}_{(B𝔡)}; fF; _{(fF )}_{(BF )}; ⟨f,B⟩𝔮; (fB)F; (f𝔲^{⋄})F]

A lax functor X [figure: F] Y is called normalized or unitary, if all 2-cells of the form A𝔡 are invertible, and a bifunctor, if is normalized and all 2-cells of the form ⟨f, g⟩𝔮 are invertible as well. Normalized lax functors and bifunctors are called strict, if the respective structural 2-cells required to be invertible are in fact identities.

dual notion: For an oplax functor X [figure: F] Y the structural 2-cells are reversed:

[figure: (fg)F; ⟨f,g⟩𝔭; (fF ) (gF ); instead of; (fF ) (gF ); ⟨f,g⟩𝔮; (fg)F]

and

[figure: AF; A𝔟; AF; instead of; AF; A𝔡; AF]

have to satisfy the duals of axioms (L0) and (L1), respectively.

1.05. Convention. Again, we use boldface Fraktur letters for the structural 2-cells. The price we have to pay for abbreviating f⟨A, B⟩F to fF is that we need to notationally distinguish the action of F on 0-cells from that of F on identity 1-cells. This can either be done by distinguishing 0-cells and identity 1-cells (e.g., A vs. 1_{A}, respectively), or by using different symbols for the functor action ( F and F , respectively), depending on the nature of the argument. Here we have opted for the second solution.

We close this section with a few remarks about the presentation of bicategorical data. Instead of the familiar pasting diagrams, we prefer string diagrams. Joyal and Street [10] attribute the first use of this technique for the manipulation of tensors to Penrose in 1971. For the author the experience of converting proofs from pasting diagrams to string diagrams was illuminating. The clear directional division of the two types of composition available in a bicategory – horizontally from left to right, and vertically from bottom to top – subjectively made the comprehension of string diagrams easier. We hope the reader will agree.

The basic entities of a bicategory are its 2-cells; they are to be thought of as labeled points. Here we have chosen to “blow up” the points to rectangular boxes with a label inside. 1-cells correspond to labeled essentially vertical lines (no horizontal tangent lines allowed), while 0-cells or objects correspond to labeled regions. Concretely, the vertical composition of two 2-cells f [figure: σ] g, and g [figure: τ] h inside a hom-category ⟨A, B⟩X and the horizontal composition of 2-cells f [figure: φ] g in ⟨A, B⟩X and h [figure: ψ] k in ⟨B, C⟩X can be depicted by either one of the first two or one of the last three diagrams, respectively:

A ^{B}_{f}^{h}^{σ}^{τ}^{g}  = ^{A} ^{B}_{f}^{h}^{σ ; τ}  and  ^{A} ^{B} ^{C}_{f} _{h}^{g} ^{k}^{φ} ^{ψ} = ^{A} ^{B}^{B} ^{C}_{f} _{h}^{g} ^{k}^{φψ}  = ^{A} ^{C}_{fh}^{gk}^{φψ}

Notice how easily string diagrams can be combined horizontally and vertically, provided the appropriate domains and codomains match. Also, since φ and ψ are independent, we may change their respective heights, which leads to φψ = fψ ; φk = φh ; gψ.

2. The general calculus of modules in a bicategory

The following crucial example was one of the reasons for Benabou [1] to introduce what we now call lax functors instead of the more restrictive concept of bifunctors.

2.00. Example. Monads may be viewed as lax functors with domain 1 , the singleton bicategory with a single element 0 that serves triple duty as a 0-cell, an identity 1-cell, and an identity 2-cell. Indeed, let 1 [figure: F] Y be a lax functor with 0F = A and write a for the endo-1-cell 0F on A. By axiom (L0) the multiplication α := ⟨0, 0⟩𝔮 from aa to a is associative, i.e.,

A ^{A} ^{A} ^{A}_{a} _{a} _{a}^{a}^{α} ^{α}^{a}  = ^{A} ^{A} ^{A} ^{A}_{a} _{a} _{a}^{a} ^{α}^{α}^{a}  (2-00)

And by axiom (L1) the 2-cell α′ := 0𝔡 from A to a serves as two-sided unit:

A ^{A} ^{A}_{A} _{a}^{a}^{α′} ^{α}^{a}  = ^{A} ^{A} ^{A}_{A} _{a}^{a}^{a𝔲_{⋄}}  and  ^{A} ^{A} ^{A}_{a} _{A}^{a} ^{α′}^{α}^{a}  = ^{A} ^{A} ^{A}_{a} _{A}^{a}^{a𝔲^{⋄}}  (2-01)

For a lax functor X [figure: F] Y each 0-cell XF in Y carries a monad structure, namely ⟨XF, ⟨X, X⟩𝔮, X𝔡⟩. We wish to construct a “bicategory of monads” over Y , through which F factors. Which 1-cells and 2-cells are appropriate? Recall that the morphisms of an ordinary category correspond to functors with domain 2 .

2.01. Definition. (0) By a morphism of monads, or m-module for short, we mean a lax functor with domain the generic 1-cell 2 ,

[figure: 0; k; 1]

(1) By an m-module homomorphism, or modism for short, we mean a lax functor with domain the generic 2-cell two

[figure: 0; m; n; ψ; 1]

Before giving a more explicit description of these concepts, let us explore a simpler notion than that of m-module, and study the composition of such entities.

2.02. Convention. From now on let Y be a fixed bicategory. To simplify the notation, we write [A, B] for the hom-category ⟨A, B⟩Y . If A′ [figure: f] A and B [figure: g] B′ are 1-cells, the functor [A, B] [figure: [f,g]] [A′, B′] pre-composes with f and post-composes with g.

2.03. Definition. Let A [figure: a] A and B [figure: b] B be endo-1-cells. A module R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ from a to b consists of a 1-cell A [figure: r] B, the carrier, equipped with a left action ar [figure: ρ_{⋄}] r and a right action rb [figure: ρ^{⋄}] r subject to the following stacking condition

A ^{A} ^{B} ^{B}_{a} _{r}  _{b}^{r}^{ρ_{⋄}} ^{ρ^{⋄}}^{r}  = ^{A} ^{A} ^{B} ^{B}_{a} _{r}  _{b}^{r} ^{ρ^{⋄}}^{ρ_{⋄}}^{r}  (2-02)

We write aR and Rb for the modules ⟨aρ_{⋄}, ar, aρ^{⋄}⟩ and ⟨ρ_{⋄}b, rb, ρ^{⋄}b⟩, respectively. If U = ⟨υ_{⋄}, u, υ^{⋄}⟩ is another module from a to b, by a modism R [figure: ψ] U we mean a 2-cell r [figure: ψ] u compatible with actions in the following sense

A ^{A} ^{B}_{a} _{r}^{u}^{ρ_{⋄}}^{ψ}^{r}  = ^{A} ^{A} ^{B}_{a} _{r}^{u} ^{ψ}^{υ_{⋄}}^{u}  and  ^{A} ^{B} ^{B}_{r}  _{b}^{u}^{ρ^{⋄}}^{ψ}^{r}  = ^{A} ^{B} ^{B}_{r}  _{b}^{u}^{ψ} ^{υ^{⋄}}^{u}  (2-03)

The category of modules from a to b with modisms as arrows is denoted by ⟨a, b⟩mod .

2.04. Proposition. If a, b and c are endo-1-cells on objects A, B and C, respectively, the composition [A, B] × [B, C] [figure: ⟨A,B,C⟩c] [A, C] restricts to a functor from ⟨a, b⟩mod × ⟨b, c⟩mod to ⟨a, c⟩mod that maps R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ and S = ⟨σ_{⋄}, s, σ^{⋄}⟩ to RS := ⟨ρ_{⋄}s, rs, rσ^{⋄}⟩. The 2-cells rbs [figure: ρ^{⋄}s] rs and rbs [figure: rσ_{⋄}] rs are modisms from RbS = (Rb)S = R(bS) to RS that constitute the ⟨R, S⟩-components of two natural transformations with codomain ⟨A, B, C⟩c .

Proof. The stacking condition for RS is trivial, since ρ_{⋄} and σ^{⋄} are independent:

A ^{A} ^{B}^{B} ^{C} ^{C}_{a} _{r}  _{s}  _{c}^{r}  ^{s}^{ρ_{⋄}s}^{rσ^{⋄}}^{r} ^{s}  = ^{A} ^{A} ^{B}^{B} ^{C} ^{C}_{a} _{r}  _{s}  _{c}^{r}  ^{s}^{ρ_{⋄}}  ^{σ^{⋄}}  = ^{A} ^{A} ^{B}^{B} ^{C} ^{C}_{a} _{r}  _{s}  _{c}^{r}  ^{s}^{rσ_{⋄}}^{ρ^{⋄}s}^{r} ^{s}

Establishing that modisms R [figure: κ] R′ and S [figure: ν] S′ yield a modism RS [figure: κν] R′S′ is similarly straightforward.

The stacking conditions for R and S, and the independence of ρ^{⋄} and σ^{⋄} as well as ρ_{⋄} and σ_{⋄} imply that ρ^{⋄}s and rσ_{⋄} are modisms from RbS to RS. Naturality turns out to be a direct consequence of the modism axioms.

The composition operations for modules defined in Proposition 2.04., although clearly associative, are not very interesting, since they ignore the right action of the first factor and the left action of the second factor. How to define a more meaningful composition operation for modules R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ and S = ⟨σ_{⋄}, s, σ^{⋄}⟩ is indicated by the second part of this proposition: form the coequalizer of ρ^{⋄}s and rσ_{⋄} in ⟨a, b⟩mod . Of course, this approach requires the existence of the appropriate coequalizers.

2.05. Definition. We say that Y has stable local colimits of a given type, provided these colimits exist in every hom-category, and 1-cell composition from either side preserves them. If Y has stable local small colimits, Carboni et. al. [3] call Y distributive.

2.06. Assumption. Y has stable local coequalizers.

For technical reasons we assume a certain choice of these coequalizers has been made. Hence we can refer to the coequalizer of a parallel pair. Although such a choice need not be canonical in any sense, we still may assume that 1-cell composition from either side preserves the chosen coequalizers.

2.07. Proposition. For endo-1-cells a and b the category ⟨a, b⟩mod has coequalizers.

Proof. For modules R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ and U = ⟨υ_{⋄}, u, υ^{⋄}⟩ from a to b consider modisms ψ and χ from R to U . Let φ be the coequalizer of ψ and χ in [A, B] . Since composition from either side preserved coequalizers, the modism axioms for ψ and χ allow us to obtain actions aw [figure: ω_{⋄}] w and wb [figure: ω^{⋄}] w via the following diagram

[figure: rb; ψb; χb; ρ^{⋄}; ub; φb; υ^{⋄}; wb; ω^{⋄}; r; ψ; χ; u; φ; w; ar; aψ; aχ; ρ_{⋄}; au; aφ; υ_{⋄}; aw; ω_{⋄}]

Since aφb is a coequalizer, and hence epi, we can derive the stacking condition for ω_{⋄} and ω^{⋄} directly from the stacking condition for υ_{⋄} and υ^{⋄} .

2.08. Definition. For modules R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ from a to b and S = ⟨σ_{⋄}, s, σ^{⋄}⟩ from b to c let ⟨φ, R • S⟩ be the coequalizer in ⟨a, b⟩mod of ρ^{⋄}s and rσ_{⋄}. We denote the carrier of R • S by r • s, and the left and right actions by ρ_{⋄} • s and r • σ^{⋄}, respectively.

2.09. Proposition. The operation • on modules extends to a family of functors

⟨a, b⟩mod × ⟨b, c⟩mod [figure: •] ⟨a, c⟩mod

that is essentially associative as specified by Axiom (B0).

Proof. To establish the functoriality of module composition, let R [figure: κ] R̄ and S [figure: ν] S̄ be modisms, where a [figure: R] b and b [figure: S] c. If φ and φ̄ are the coequalizers used to construct R • S and R̄ • S̄, respectively, the universal property of φ induces κ • ν via

[figure: RbS; ρ^{⋄}s; rσ_{⋄}; κbν; RS; φ; κν; R • S; κ • ν; _{R̄bS̄}; _{_{ρ̄^{⋄}}S̄}; _{R̄σ̄_{⋄}}; _{R̄S̄}; φ̄; _{R̄ • S̄}]

where the rectangles on the left commute because of naturality, cf. Proposition 2.04..

For another module U = ⟨υ_{⋄}, u, υ^{⋄}⟩ from c to d and the coequalizers φ, λ, ψ and μ used in the construction of R•S, S •U , (R•S)•U and R•(S •U) , respectively, consider the diagram

[figure: Rb(S • U); _{ρ^{⋄}(s • u)}; r(σ_{⋄} • u); RbSU; _{ρ^{⋄}su}; rσ_{⋄}u; rbλ; R(S • U); μ; R • (S • U); _{⟨R,S,U⟩𝔞̇}; RbScU; _{ρ^{⋄}scu}; rσ_{⋄}cu; rbσ^{⋄}u; rbsυ_{⋄}; RSU; φu; rλ; RScU; φcu; rσ^{⋄}u; rsυ_{⋄}; (R • S)U; ψ; (R • S) • U; _{(R • S)cU}; (r • σ^{⋄})u; _{(r • s)υ_{⋄}}]  (2-04)

Both φu ; ψ and rλ ; μ turn out to be multi-coequalizers of the four 2-cells ρ^{⋄}σ^{⋄}u, ρ^{⋄}sυ_{⋄}, rσ_{⋄}υ_{⋄}, and rσ_{⋄}cu ; rσ^{⋄}u = rbσ^{⋄}u ; rσ_{⋄}u from RbScU to RSU , which induces the iso ⟨R, S, U⟩𝔞̇ from (R • S) • U to R • (S • U) . The universal property of colimits also allows us to establish the coherence axiom (B0) in a lengthy but straightforward diagram chase.

From now on we will no longer explicitly mention the associativity isos. In some sense it is taken care of by the geometry of the diagrams, this applies to pasting diagrams as well as to string diagrams. More formally, we utilize the fact that every bicategory is biequivalent to a 2-category, e.g., Theorem 1.4 in [4].

3. Interpolads and monads

In general, endo-1-cells, modules and modisms do not form a bicategory, since identity modules may not exist. Identity modules must at least be idempotent under module composition. The only reasonable candidate for an identity module on A [figure: a] A is a, equipped with suitable left and right actions. Of course, modules now have to be compatible with the extra structure.

3.00. Definition. Let A [figure: a] A be a 1-cell in Y . We refer to any 2-cell aa [figure: α] a as a multiplication, and we call α interpolative, and the pair A = ⟨a, α⟩ an interpolad, if α is a coequalizer of αa and aα (and hence in particular associative). If A and B = ⟨b, β⟩ are interpolads, an i-module R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ from A to B, denoted by A [figure: R] B, is a module from a to b that satisfies

(I0) ρ_{⋄} is a coequalizer of αr and aρ_{⋄} (not necessarily the chosen one);

(I1) ρ^{⋄} is a coequalizer of ρ^{⋄}b and rβ (not necessarily the chosen one).

Modisms between i-modules are ordinary modisms.

It is clear that interpolads can be interpreted as their own identity i-modules, with the multiplication serving double duty as left and right action; the two i-module axioms together with the stacking condition (2-02) guarantee the properties of a left and right unit, respectively.

3.01. Remarks. (0) Considering idempotent modules ⟨α_{⋄}, a, α^{⋄}⟩ instead of interpolads leads to no further generality. In this case for an i-module R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ from ⟨α_{⋄}, a, α^{⋄}⟩ to ⟨β_{⋄}, b, β^{⋄}⟩ the left action ρ_{⋄} ought to be a coequalizer of α^{⋄}r and aρ_{⋄}, and the right action ρ^{⋄} ought to be a coequalizer of ρ^{⋄}b and rβ_{⋄}. For ⟨α_{⋄}, a, α^{⋄}⟩ to be an endo-i-module on itself, both α_{⋄} and α^{⋄} have to be coequalizers of α^{⋄}a and aα_{⋄}.

(1) Another way to arrive at the notions of interpolad and i-module is to relax the notions of lax functor with domain 1 , respectively 2 , just enough as to keep the coequalizers previously guaranteed by implicit split coequalizers. [more precisely...]

(2) Each Y -object A trivially is structured as an interpolad: with the identity 1-cell A as carrier and the multiplication A𝔲 (cf. Remark 1.02.(0)). A 1-cell A [figure: r] B can be viewed as a trivial i-module between trivial interpolads, by taking Ar [figure: r𝔲_{⋄}] r and rB [figure: r𝔲^{⋄}] r as actions. Naturality and coherence (use Remark 1.02.(1) here) guarantee that the diagram (2-02) commutes, and also the coequalizer conditions. Clearly, 2-cells of Y turn into modisms between such trivial i-modules.

(3) The terms “interpolative” and “interpolad” are inspired by the interpolation property for binary relations that is of some interest in domain theory. A relation a ⊆ A × A has the interpolation property, if it satisfies a ⊆ aa, i.e., the converse of the transitivity requirement. The latter, in general, corresponds to the associativity of the multiplication, while the interpolation property may be viewed as a couniversal property of the associativity diagram. We propose the slogan:

associative + interpolation property = interpolative

3.02. Theorem. Interpolads in Y , i-modules and modisms form a bicategory Y -int that contains Y as a full sub-bicategory.

Proof. I-modules between two interpolads and their modisms clearly form a category. A simple coequalizer argument shows that i-modules are closed under module composition.

For an i-module A [figure: R] B let ar [figure: φ] a • r be the coequalizer of αr and aρ_{⋄}. Since ρ_{⋄} is a coequalizer of this pair as well, we obtain an iso A • R [figure: R𝔲̇_{⋄}] R. Similarly, we get an iso R • B [figure: R𝔲̇^{⋄}] R. These isos together with the associativity isos 𝔞̇ defined in the proof of Proposition 2.09. clearly satisfy coherence axiom (B1). This establishes the bicategory Y -int .

Naturality of 𝔲_{⋄} and (I0) for an i-module ⟨ρ_{⋄}, r, ρ^{⋄}⟩ from ⟨A, A𝔲⟩ to ⟨B, B𝔲⟩ imply

A ^{A} ^{A} ^{B}_{A} _{A} _{r}^{r} ^{ρ_{⋄}}^{r𝔲_{⋄}}^{r}  = ^{A} ^{A} ^{B} ^{B}_{A} _{A} _{r}^{r}^{A𝔲}^{ρ_{⋄}}^{A}  = ^{A} ^{A} ^{A} ^{B}_{A} _{A} _{r}^{r} ^{ρ_{⋄}}^{ρ_{⋄}}^{r}

Since Aρ_{⋄} is epi, we get ρ_{⋄} = r𝔲_{⋄}. A similar argument shows ρ^{⋄} = r𝔲^{⋄}. Hence we can identify modules from A𝔲 to B𝔲 with 1-cells from A to B. Coherence condition (B1) implies that composition of trivial i-modules reduces to 1-cell composition in Y .

3.03. Proposition. If A and B are interpolads with carriers a and b, respectively, and if a [figure: R] b is any module, then A • R • B is an i-module from A to B.

Proof. Let R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ be a module from a to b, and consider the following diagram, where φ is the chosen coequalizer of αr and aρ_{⋄} used in the construction of a • r:

[figure: aa(a • r); α(a • r); a(α • r); aaar; αar; aαr; aaφ; a(a • r); α • r; aaaar; aαar; aaαr; aaαr; aaaρ_{⋄}; aar; αr; aφ; a • r; aaar; aαr; aαr; aaρ_{⋄}; ar; φ; aar; αr; aρ_{⋄}]

Since 1-cell composition preserves coequalizers, and since α is a coequalizer of αa and aα, it follows immediately that the left action α • r of A • R is a coequalizer of α(a • r) and a(α • r) and hence satisfies condition (I0). Similarly, we see that the left action α • r • b of A • R • B satisfies (I0), and that the right action a • r • β of this module satisfies (I1), i.e., A • R • B is an i-module from A to B.

Clearly, an interpolative multiplication is the minimal structure necessary to support identity modules, and hence yield a bicategory. Adding more structure will lead to sub-bicategories of Y -int that in general will not be full, if the modules have to preserve this additional structure. The modisms will be the same in all cases.

The obvious extra structure to add is a two-sided unit for the multiplication. Then we obtain monads instead of interpolads. The interpolation property for the multiplication then follows from the unit axioms, hence only associativity has to be required.

3.04. Proposition. Let A = ⟨a, α, α′⟩ and B = ⟨b, β, β′⟩ be monads with units α′ and β′, respectively. For a module R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ from a to b the following are equivalent:

(a) R is an m-module;

(b) R is a module that is compatible with the multiplication

A ^{A} ^{B} ^{B}_{a} _{a} _{r}^{r}^{α} ^{ρ_{⋄}}^{a}  = ^{A} ^{A} ^{A} ^{B}_{a} _{a} _{r}^{r} ^{ρ_{⋄}}^{ρ_{⋄}}^{r}  ;  ^{A} ^{B} ^{B} ^{B}_{r}  _{b}  _{b}^{r} ^{β}^{ρ^{⋄}}^{b}  = ^{A} ^{B} ^{B} ^{B}_{r}  _{b}  _{b}^{r}^{ρ^{⋄}} ^{ρ^{⋄}}^{r}

(3-00)

and with the units

A ^{A} ^{B}_{A} _{r}^{r}^{α′} ^{ρ_{⋄}}^{a}  = ^{A} ^{A} ^{B}_{A} _{r}^{r}^{r𝔲_{⋄}}  ;  ^{A} ^{B} ^{B}_{r} _{B}^{r} ^{β′}^{ρ^{⋄}}^{b}  = ^{A} ^{B} ^{B}_{r} _{B}^{r}^{r𝔲^{⋄}}  (3-01)

(c) R is an i-module, and both actions are split epi.

Proof. (b) ⇔ (a) Let F be a lax functor from 2 to Y . Besides two monads A and B (based on 0F and 1F , respectibely) F also specifies a module R based on kF subject to five compatibility conditions: the stacking condition (2-02) as well as the commutativity of the diagrams specified by (I0) and (I1), and the two equalities in (3-01).

(b) ⇒ (c) By (3-01) the actions are split epi, which togther with (3-00) implies (I0) and (I1).

(c) ⇒ (b) Let r [figure: λ] ar be a left inverse for ρ_{⋄}. Then we have

A ^{A} ^{B}_{A} _{r}^{r}^{α′} ^{ρ_{⋄}}^{a}  = ^{A} ^{A} ^{B}_{A} _{r}^{r}^{α′}  ^{λ}^{ρ_{⋄}}^{ρ_{⋄}}^{a} ^{a}^{r}^{r}  = [figure: A; A; B; A; A; r; r; α′; λ; α; ρ_{⋄}; a; a; r; a] = ^{A} ^{A} ^{B}^{A}_{A} _{r}^{r} ^{λ}^{a𝔲_{⋄}}^{ρ_{⋄}}^{a} ^{r}^{a}  = ^{A} ^{A} ^{B}_{A} _{r}^{r}^{r𝔲_{⋄}}^{λ}^{ρ_{⋄}}^{r}^{a}^{r}

where the second step uses (I0), the third one uses (2-01), and the fourth one uses the naturality of 𝔲_{⋄}. The other condition follows by symmetry.

3.05. Remarks. (0) Condition (b) of Proposition 3.04. directly translates the notion of bimodule familiar from algebra, cf. Examples 1.03.(2), 3.07.(1) as well as [7] and [3]. We have dropped the prefix “bi”, because of it’s different connotation in “bicategory”. The terms “distributor” and “profunctor” also appear in the literature.

(1) Remark 3.01.(2) carries over to the monad setting.

3.06. Theorem. Monads in Y , m-modules and modisms form a bicategory Y -mnd that contains Y as a full sub-bicategory. Moreover, there is a forgetful lax functor Y -mnd [figure: 𝔜] Y with the property that every lax functor X [figure: F] Y factors through 𝔜 by means of a normalized lax functor.

Proof. A simple calculation shows that m-modules are closed under module composition. Thus we obtain a bicategory Y -mnd . The lax functor 𝔜 maps monads to their underlying objects, m-modules to their underlying 1-cells, and modisms to their underlying 2-cells. The structural 2-cells for composits are given by the coequalizers used to define the m-module composition, and by the units of the monads. Hence by construction any lax functor X [figure: F] Y factors through 𝔜 in the specified fashion.

In general, a lax functor X [figure: F] Y will not factor through 𝔜 by means of a bifunctor. The price for upgrading a lax functor to a normalized lax functor is that 1-cell composition in Y -mnd usually is more complicated than in Y .

3.07. Examples. (0) Let X be the suspension of set (or of any elementary topos), i.e., X has a single 0-cell ∗, sets as (endo-) 1-cells with cartesian product as horizontal composition, and functions as 2-cells. Here we use capital letters for sets, and small letters for elements. A module R from A to B is a set R equipped with left and right actions A × R [figure: ρ_{⋄}] R and R × B [figure: ρ^{⋄}] R that satisfy

(ar)b = a(rb)  (3-02)

for a ∈ A, r ∈ R, and b ∈ B, where we have used concatenation for both actions. Composing R with a module S from B to C results in the module R • S based on the set (R × S)/ ↭, where ↭ is the smallest equivalence relation on R × S that contains all pairs ⟨⟨rb, s⟩, ⟨r, bs⟩⟩, for r ∈ R, b ∈ B, and s ∈ S. Modisms are simply functions that commute with the actions.

Interpolads now are interpolative semi-groups, i.e., semi-groups ⟨A, ∘⟩ with surjective multiplication such that the kernel pair of ∘ coincides with the equivalence relation generated by ⟨∘×A, A×∘⟩ (this automatically implies associativity). The actions of an i-module from ⟨A, ∘⟩ to an interpolative semi-group ⟨B, ·⟩ besides (3-02) have to satisfy

(a′ ∘ a)r = a′(ar) and (rb)b′ = r(b · b′)  (3-03)

Moreover, they have to be surjective, and their kernel pairs have to agree with the equivalence relations generated by ⟨∘×R, A×ρ_{⋄}⟩ and ⟨ρ^{⋄} ×B, R×·⟩, respectively.

Semi-groups with identity, i.e., monoids, automatically are interpolative and correspond to monads. The actions of an m-module from a monoid ⟨A, ∘, e⟩ to a monoid ⟨B, ·, i⟩ besides satisfying (3-02) and (3-03) have to preserve the identities, i.e.,

er = r and ri = r  (3-04)

which again implies the interpolation properties.

If ⟨ρ_{⋄}, R, ρ^{⋄}⟩ is an m-module from ⟨A, ∘, e⟩ to ⟨B, ·, i⟩, then invertible elements x ∈ A and y ∈ B give rise to i-modules between these monoids that need not be mmodules: just defined new left and right actions A × R [figure: ρ̄_{⋄}] R and R × B [figure: ρ̄^{⋄}] R by ⟨a, r⟩ρ̄_{⋄} := (x ∘ a)r and ⟨r, b⟩ρ̄^{⋄} := r(b · y) . I-modules arising in this fashion might be called affine.

For interpolative semi-groups the multiplication is a left and a right action and induces an idempotent endo-module, the identity module. For ordinary semi-groups this can fail. In case of the positive natural numbers N_{+} under addition we obtain N_{+} • N_{+} ≅ N_{>1} , on which N_{+} acts from both sides.

Just as monoids are categories with one object, m-modules between monoids can be interpreted as special categories with two objects, cf. part (4) below.

(1) Replacing set above with ab , the category of abelian groups, yields unitary rings as monads. Bimodules and their homomorphisms serve as m-modules and modisms. The composition of bimodules is their tensor product, cf. Example 1.03.(2).

For the suspension of ab , interpolads will be “rngs”, i.e., “rings not necessarily with unit”, that satisfy the interpolation property. Since regular epis in ab are cokernels and hence surjective, the same description as in (0) applies.

(2) For Y = rel , the bicategory of sets, relations and inclusions, an endo-1-cell on A is just a binary relation a ⊆ A × A. A module from a to b ⊆ B × B is a relation r ⊆ A × B that satisfies ar ⊆ r ⊇ rb. Since coequalizers in rel are trivial, modules compose like ordinary relations. Modisms are just inclusions.

The interpolads in rel turn out to be Steve Vickers’ information systems [12], i.e., sets equipped with an idempotent (= transitive and interpolative) relation. An imodule from ⟨A, <⟩ to ⟨B, ⊏⟩ now is a relation r ⊆ A × B that satisfies ( < r) = r = (r ⊏ ) . Since the relation < on 2 is transitive but not idempotent, we see again that identity modules may fail to exist in the absence of the interpolation property. The resulting bicategory, which we call kar , also has been studied by Rosebrugh and Wood [8]. However, their motivation was not a weakening of the notion of monad, but rather the issue of Cauchy-completeness. We return to this subject in Section 4.

A monad in rel is simply a pre-ordered set i.e., a set equipped with a reflexive and transitive relation (sometimes called graph). M-modules from ⟨A, ≤⟩ to ⟨B, ⊑⟩ are those relations r ⊆ A × B that satisfy ( ≤ r) ⊆ r ⊇ (r ⊑ ) (by reflexivity the other inclusions hold automatically) or, equivalently, for which the characteristic function preserves order from ⟨A, ≤⟩^{op} × ⟨B, ⊑⟩ to 2 . Such relations are known as order-ideals. Pre-ordered sets, order-ideals and inclusions form a sub-bicategory idl of kar . Since kar is locally ordered, this inclusions happens to be full.

(3) Every ord -enriched category Y may be viewed as a bicategory – the homcategories are ordered sets and the composition preserves order. Coequalizers in the hom-categories are identities (or isos, if we do not insist on antisymmetry) and hence are trivially preserved. Endo-1-cells on A are elements of the poset [A, A] , and a module from a ∈ [A, A] to b ∈ [B, B] is an element r ∈ [A, B] that satisfies ar ≤ r ≥ rb. In this situation we need not distinguish between i-modules and m-modules: the actions simply have to be identities, i.e., we have ar = r = rb. For Y = ord interpolads correspond to idempotent endo-functions, while monads correspond to closure operators. Another special case is the bicategory rel , cf. part (2). It is even enriched in the category cslat of complete lattices with additive or sup-preserving (= left adjoint) functions, since relations from A to B bijectively correspond to left-adjoint functions from AP to BP . Categories enriched in cslat are also known as quantaloids, cf. Example 5.11.(2).

(4) Just as a relation between sets A and B can be viewed as a 2-valued A×B matrix, a span A [figure: r]B, i.e., a function R [figure: r] A × B, can be viewed as a set -valued A × B matrix, assigning to each pair ⟨x, y⟩ ∈ A × B a set of formal arrows. The composition with B [figure: s]C now is just the matrix product with cartesian product × and disjoint union + instead of boolean meet ∧ and join ∨, i.e., ⟨x, z⟩(rs) is given by ∑_{y∈B} ⟨x, y⟩r × ⟨y, z⟩s. The identity matrices for this operation have singletons in the diagonal, and the empty set everywhere else. Span-morphisms between spans from A to B are A × B matrices of functions. Their vertical composition as well as the formation of (co-)limits is performed component-wise. The resulting bicategory is called spn .

An endo-span a on A can be interpreted as a directed multigraph with vertex set A. If b is another such with vertex set B, a module a [figure: R] b specifies sets of formal arrows from A-vertices to B -vertices that compose associatively with the a-arrows and the b-arrows. If b [figure: S] c is another module, the composite RS at each B -vertex simply links incoming and outgoing arrows, while R • S takes the possible 1-step transfers within the directed multigraph b into account, and factors by the appropriate equivalence relation.

Two pairs of arrows ⟨f, h̄⟩ ∈ ⟨x, u⟩r × ⟨u, z⟩s and ⟨f̄, h⟩ ∈ ⟨x, v⟩r × ⟨v, z⟩s have a common pre-image wrt. ρ^{⋄}s and rσ_{⋄} iff ⟨f, h̄⟩ “diagonalizes” over ⟨f̄, h⟩, i.e., iff f̄ = f ; g and h̄ = g ; h for some g ∈ ⟨u, v⟩b, or vice versa. Hence ⟨f, h̄⟩ and ⟨f̄, h⟩ are equivalent iff they can be linked by a finite “zig-zag” of diagonals. The corresponding relation ↭_{⟨x,z⟩} is defined on ∑_{y∈B} ⟨x, y⟩r×⟨y, z⟩s, the ⟨x, z⟩- component of rs. Just as span-morphisms are matrices of functions, here we have a matrix ↭ of relations, a span-relation on rs.

A span morphism aa [figure: α] a equips a directed multigraph a with a composition operation, i.e., a family of functions ∑_{y∈A} ⟨x, y⟩a × ⟨y, z⟩a [figure: ⟨x,z⟩α] ⟨x, z⟩a. Clearly, associativity of α is equivalent to the associativity of the composition operation. If two vertices are connected by two different directed paths, we may ask if one path can be “deformed” into the other, using the composition operation. Geometrically speaking, this is possible iff the “region” between the paths can be completely “triangulated”. Thus we obtain a notion of homotopy equivalence between paths. The pair ⟨a, α⟩ now is an interpolad provided that for paths of length 2 with common domain and codomain the homotopy equivalence (the same as ↭ above) is the kernel of the composition operation.

It is now obvious that a monad ⟨a, α, α′⟩ in spn on A turns out to be a small category. The ⟨x, x⟩-component 1 [figure] ⟨x, x⟩a of α′ selects the identity morphism on x ∈ A. M-modules from A to another monad B correspond to profunctors, i.e., functors \xymatrix{A^{op} × B \ar[r] & set} . Small categories, profunctors and appropriate natural transformations form a bicategory prof . In fact, an m-module from A to B may be viewed as a special category on the disjoint union A + B that contains A and B as full subcategories and has no arrows from B-objects to A-objects.

Paré and Wood briefly considered the notion of “category without identities” in an attempt to generalize idempotent relations. For their so-called taxonomies they just dropped the requirement for identities. But without interpolative composition, identity modules, i.e., generalized hom-functors, need not exist. Hence we propose interpolads in spn as more useful taxonomies. The following graphs on {w, x, y, z} illustrate the difference. In the first one the interpolation property fails since both triangles commute, but there is no “zig-zag” linking y and w. The central trapezoid of the second graph does not commute, hence the interpolation property is satisfied.

[figure: y; x; z; w; y; x; ≠; z; w]

Hayashi [5] introduced the notion of semi-functor between categories by dropping the requirement for functors that identities have to be preserved, cf. also [6]. Any semi-functor between (small) categories that is not a functor induces a module between monads in spn that is an i-module, but not an m-module.

(5) The construction in (4) works for any complete and cocomplete monoidal category V . The bicategory V -spn has sets as objects and matrices of V -objects as 1-cells. Monads in V -spn then correspond to small V -categories.

Already Benabou [1] introduced a “multi-object” version of the notion of monad in Y , which he termed “polyad”. Later this was interpreted as a “category enriched in Y ”, cf., e.g., [9] and the historical remarks therein. It is straightforward to give a similar generalization of interpolads, resulting in the notion of “taxonomy enriched in a bicategory”. Instead of coequalizers, more complicated colimits are needed. This also applies to the corresponding notion of i-module. We leave it to the reader to spell out the details.

4. Splitting idempotents

Recall that an idempotent morphism f in a category is said to split, if it can be factored as a split epi e followed by a split mono m such that m ; e is an identity. A category is called Cauchy-complete, if all idempotents split. A Cauchy-completion of a category C can, e.g., be realized within C /C by taking all idempotent morphisms as objects and only those morphisms, where both components coincide. These notions immediately carry over to locally ordered bicategories. The failure of rel to be Cauchy-complete in this sense prompted Rosebrugh and Wood [8] to work with kar instead, which happens to be the Cauchy-completion of rel , cf. Example 3.07.(2).

In a general bicategory already the notion of idempotent 1-cell becomes more subtle. Requiring aa = a for a 1-cell A [figure: a] A certainly is too strong, we rather should specify an iso aa [figure: α] a. But in order to be able to replace any power a^{n}, n > 0 , by a, which seems to capture the spirit of idempotency, we must ask α to induce only one iso from each such power to a, i.e., to be associative. As an iso, α then automatically is a coequalizer of aα and αa, i.e., we are looking at special interpolads.

4.00. Definition. (0) An idempolad ⟨a, α⟩ on A ∈ \underline{Y} consists of a 1-cell A [figure: a] A and an iso aa [figure: α] a that satisfies aα = αa.

(1) A bicategory Y is called Cauchy-complete, if all idempolads in Y split.

(2) The Cauchy-completion Y -chy of a bicategory Y is the smallest Cauchy-com-plete bicategory that contains Y as a full sub-bicategory.

4.01. Examples. (0) The interpretation of Y -objects as trivial interpolads does in fact yield idempolads.

(1) If Y is the suspension of a cartesian category C , then for every idempolad ⟨X, x⟩ in C Σ the object X is sub-terminal in C : consider C -arrows f, g ∈ ⟨C, X⟩C ; since the bijections X × x and x × X coincide, the projections from X × X to X agree on ⟨f, f, f⟩ ; (X × x) and ⟨f, g, f⟩ ; (x × X) . Therefore ⟨f, f, f⟩ = ⟨f, g, f⟩, and hence f = g.

In particular, for C = set we see that although every infinite set X is isomorphic to its square, no isomorphism from X × X to X can be associative. Hence the only idempolads in the suspension of set are singletons, which trivially split.

4.02. Proposition. Y -int is Cauchy-complete.

Proof. Consider an endo-i-module R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ on an interpolad ⟨a, α⟩ and an associative isomorphism R • R [figure: μ] R. Let φ, φ′, and φ′′ be the coequalizers used in the construction of R • R, (R • R) • R, and R • (R • R) , respectively, cf. Definition 2.08.. Since both φR ; φ′ and Rφ ; φ′′ are multi-coequalizers of the four 2-cells ρ^{⋄}ρ^{⋄}r, ρ^{⋄}rρ_{⋄}, rρ_{⋄}ρ_{⋄} and rρ_{⋄}ar ; rρ^{⋄}r = raρ^{⋄}r ; rρ_{⋄}r from RaRaR to RRR, (cf. the proof of Proposition 2.09., and in particular Diagram (2-04)), we now have

[figure: RR; φ; (R • R)R; φ′; μR; R • R; μ; RRR; Rφ; φR; R • R • R; R • μ; μ • R; R; R(R • R); Rμ; φ′′; R • R; μ; RR; φ]  (4-00)

Therefore (φ ; μ)R ; φ = R(φ ; μ) ; φ has the same multiple coequalizer property, and it follows immediately that φ, and hence also φ ; μ is in fact a coequalizer of (φ ; μ)R and R(φ ; μ) . Hence ⟨r, φ ; μ⟩ is an interpolad in Y .

To see that U := ⟨ρ_{⋄}, r, φ ; μ⟩ and V := ⟨φ ; μ, r, ρ^{⋄}⟩ are i-modules from ⟨a, α⟩ to ⟨r, φ ; μ⟩ and back, observe that in both cases the stacking condition follows directly from φ ; μ being modism. Axioms (I0) and (I1) are satisfied, since R is an i-module and since ⟨r, φ ; μ⟩ is an interpolad. Finally, it is easy to see that U • V = R • R ≅ R, while V • U essentially is the identity module on ⟨r, φ ; μ⟩.

We expect the full sub-bicategory Y -idm of Y -int , spanned by the idempolads, to be the Cauchy-completion of Y . Since the notion of idempolad is self-dual, in case that Y locally has stable equalizers as well, we would also expect to recover the Cauchycompletion of Y as a full sub-bicategory inside the bicategory of co-interpolads. Hence i-modules between idempolads ought to be i-comodules and their composition must be independent of the local (co-)completeness properties of Y .

4.03. Proposition. I-modules between idempolads have isomorphisms as actions and bijectively correspond to i-comodules. Moreover, if B = ⟨b, β⟩ is an idempolad, then any i-modules A [figure: R] B [figure: S] C satisfy R • S = RS.

Proof. Set R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩. Since ρ^{⋄} is a coequalizer of ρ^{⋄}b and rβ, stability implies that ρ^{⋄}b is a coequalizer of ρ^{⋄}bb and rβb = rbβ. In particular ρ^{⋄}bb ; ρ^{⋄}b = rbβ ; ρ^{⋄}b = ρ^{⋄}bb ; rβ. But ρ^{⋄}bb is epi, so we obtain ρ^{⋄}b = rβ and therefore ρ^{⋄} is an isomorphism. If A is an idempolad as well, we see that ρ_{⋄} is also an isomorphism, and clearly ⟨(ρ_{⋄})^{−1}, r, (ρ^{⋄})^{−1}⟩ is an i-comodule.

For the second i-module S = ⟨σ_{⋄}, s, σ^{⋄}⟩ we get βs = bσ_{⋄}, and therefore ρ^{⋄}bs = rbσ_{⋄}. Now ρ^{⋄}bs ; rσ_{⋄} = rbσ_{⋄} ; ρ^{⋄}s implies rσ_{⋄} = ρ^{⋄}s. Hence the coequalizer φ used in the construction of r • s is an isomorphism.

4.04. Proposition. For every bicategory X the bicategory X -idm of idempolads, i-modules and modisms is a Cauchy-completion of Y .

Proof. As we just saw in Proposition 4.03., the composition of i-modules between idempolads is well-defined for any bicategory. This establishes the bicategory X -idm . Its Cauchy-completeness follows just as that of Y -int in Proposition 4.02. and the minimality is obvious.

What about the Cauchy-completion of Y -mnd ? Under mild hypotheses on the hom-categories every interpolad in Y occurs by splitting an idempolad in Y -mnd .

4.05. Proposition. If Y is distributive, for an interpolad ⟨a, α⟩ the 1-cell A [figure: a+A] A carries a monad structure with multiplication [α, a𝔲_{⋄}, a𝔲^{⋄}] + A𝔲 and unit A [figure: Aι] a + A. Moreover, ⟨α, a, α⟩ is an idempotent m-module on this monad, which in Y -int splits through ⟨a, α⟩.

Proof. By distributivity, (a + A)(a + A) is isomorphic to aa + aA + Aa + AA. Now the 2-cells

aa [figure: α] a , Aa [figure: a𝔲_{⋄}] a , aA [figure: a𝔲^{⋄}] a and AA [figure: A𝔲] A

induce a 2-cell β := [α, a𝔲_{⋄}, a𝔲^{⋄}] + A𝔲 from (a + A)(a + A) to a + A that is easily seen to be associative. The coprojection A [figure: Aι] a + A serves as a 2-sided unit. Hence ⟨a + A, β, Aι⟩ is a monad.

On a we now define a left action (a + A)a [figure: ρ_{⋄}] a via α and a𝔲_{⋄}, and a right action a(a + A) [figure: ρ^{⋄}] a via α and a𝔲^{⋄}. A straightforward calculation establishes the m-module axioms, and the fact that this m-module in Y -int splits through the original interpolad ⟨a, α⟩.

The Cauchy-completion of Y -mnd has idempotent m-modules as objects. These may be viewed as pairs consisting of a monad together with a compatible interpolad. The bifunctor to Y -int that forgets the monad is unlikely to be full, unless each interpolad can only occur with essentially one monad. This does happen in the locally ordered case, e.g., for rel : an idempotent relation on a set A can be an idempotent m-module only for the preorder obtaind by forming the union with the diagonal on A. In case of spn this fails, since forming disjoint unions with identity modules is not an idempotent operation. Hence in general we cannot realize the Cauchy-completion of Y -mnd within Y -int .

5. The inheritance of closedness

The familiar notion of closedness for a symmetric monoidal category splits into two notions in the absence of symmetry: left-closedness and right-closedness. Their usefulness in connection with m-modules was already observed by Lawvere [7]. These notions make perfect sense in any bicategory. Street and Walters [11] have formalized the corresponding generalizations in terms of the existence of all right extensions respectively all right liftings:

5.00. Definition. For 1-cells A [figure: r] B and A [figure: t] C, a right extension along r of t consists of a 1-cell B [figure: r▷t] C and a 2-cell r(r ▷ t) [figure: ev] t such that for every s ∈ [B, C] pasting at r▷t of ev is bijective from ⟨s, r ▷ t⟩[B, C] to ⟨rs, t⟩[A, C] , indicated by dashed 2-cells in the following diagrams:

[figure: A; B; C; r; s; t; φ; ev; r▷t; =; A; B; C; r; s; t; ψ]

dual notion: Right liftings (morally: left extensions) in Y are right extensions in Y^{op} ; they are denoted by (t ◁ s)s [figure: ve] s.

There is an unfortunate name-mismatch between the terms “left-closed” and “rightclosed” on one side, and the terms “right extension” and “right lifting” on the other side.

In analogy to the monoidal case, we talk about (exponential) transposition and currying when referring to the operation of pasting at a right extension or a right lifting and its inverse. These are frequently expressed by means of a notation borrowed from natural deduction:

[figure: s; r ▷ t; rs; t; r; t ◁ s]

The existence of all right extensions, resp. all right liftings can be expressed equivalently as follows: for any Y -objects A, B, and C, there exist adjunctions

[figure: [B, C]; r▷−; [r,C]; ⊤; [A, C]; resp.; [A, B]; −◁s; [A,s]; ⊤; [A, C]]  (5-00)

Consequently, 1-cell composition from the left resp. from the right is left-adjoint and hence preserves colimits, in particular coequalizers. We always write ev for the counit of the first adjunction (reminding us of “evaluation”), and ve for the counit of the second adjunction, regardless of which 2-cells r and s, respectively, are involved. This, we hope, will keep the notation manageable. In particular, r(r ▷ t) [figure: ev] t is a transpose of id_{r▷t}, and (t ◁ s)s [figure: ve] t is a transpose of id_{t◁s}. Because of the obvious duality, we will restrict our attention to the construction of right extensions. In addition to the existence of stable local coequalizers in Y (cf. Assumption 2.06.), for the rest of this section we require

5.01. Assumption. Y has all right extensions.

Just as r▷t can be formed whenever r and t have the same domain, the expression φ ▷ χ for 2-cells φ and χ is meaningful whenever the horizontal domains agree. This generality will greatly simplify computation with right extensions.

5.02. Definition. For 2-cells q [figure: φ] r and tu [figure: χ] v with A [figure: q,r] B, A [figure: t] C [figure: u] D and A [figure: v] D, we define the 2-cell φ ▷_{u} χ by means of:

[figure: A; B; C; D; q; r▷t; u; v; φ; ev; χ; r; t; =; A; B; C; D; q; r▷t; u; v; φ▷_{u}χ; ev; q▷v]  (5-01)

If u happens to be an identity 1-cell, we drop the subscript.

If all right liftings exist, φ▷_{u}χ decomposes as (φ ▷ χ^{◁})u ; ve , where χ^{◁} is a curried form of χ. But the following arguments do not need right liftings.

5.03. Proposition. For a module T = ⟨τ_{⋄}, t, τ^{⋄}⟩ from a to c let τ_{⋄}^{▷} from t to a ▷ t be the curried form of τ_{⋄}. The 1-cell a [figure: a▷t] c together with the transpose of a ▷ τ_{⋄}^{▷} as left action and a ▷_{c} τ^{⋄} as right action constitutes a module a ▷ T from a to c.

Proof. The stacking condition for the specified actions follows from

A ^{A} ^{A} ^{C} ^{C}_{a} _{a} _{a▷t} _{c}^{t}^{a▷τ_{⋄}^{▷}}^{ev}^{a▷_{c}τ^{⋄}}^{ev}^{a▷(a▷t)}^{a▷t}^{a▷t}  = ^{A} ^{A} ^{A} ^{C} ^{C}_{a} _{a} _{a▷t} _{c}^{t}^{ev}^{τ_{⋄}} ^{τ^{⋄}}^{t}^{t}  = ^{A} ^{A} ^{A} ^{C} ^{C}_{a} _{a} _{a▷t} _{c}^{t}^{ev} ^{τ^{⋄}}^{τ_{⋄}}^{t}^{t}  = ^{A} ^{A} ^{A} ^{C} ^{C}_{a} _{a} _{a▷t} _{c}^{t} ^{a▷_{c}τ^{⋄}}^{a▷τ_{⋄}^{▷}}^{ev}^{ev} ^{a▷t}^{a▷(a▷t)}^{a▷t}

since the right branches under a(a ▷ t) [figure: ev] t in the outer diagrams have to agree.

5.04. Proposition. If a, b and c are endo-1-cells on objects A, B and C, respectively, the functor [A, B]^{op} × [A, C] [figure: ▷] [B, C] restricts to a functor from (⟨a, b⟩mod )^{op}× ⟨a, c⟩mod to ⟨b, c⟩mod that maps R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ and T = ⟨τ_{⋄}, t, τ^{⋄}⟩ to the module R ▷ T with carrier r ▷ t, left action the transpose of ρ^{⋄} ▷ t, and right action r ▷_{c} τ^{⋄}. Furthermore, if τ_{⋄}^{▷} denotes the curried version of τ_{⋄}, the 2-cells r ▷ t [figure: ρ_{⋄}▷t] ar ▷ t and r ▷ t [figure: r▷τ_{⋄}^{▷}] r ▷ (a ▷ t) = ar ▷ t are modisms from R ▷ T to aR ▷ T = R ▷ (a ▷ T) that constitute the ⟨R, T ⟩-components of two natural transformations with domain ▷.

Proof. The stacking condition for b(ρ_{⋄} ▷ t); ev and r ▷ τ_{⋄}^{▷} follows from

A ^{B} ^{B} ^{C} ^{C}_{r}  _{b} _{r▷t} _{c}^{t}^{ρ^{⋄}▷t}^{ev}^{r▷_{c}τ^{⋄}}^{ev}^{b▷(r▷t)}^{r▷t}^{r▷t}  = ^{A} ^{B} ^{B} ^{C} ^{C}_{r}  _{b} _{r▷t} _{c}^{t}^{ρ^{⋄}} ^{ev} ^{τ^{⋄}}^{t} ^{t}  = ^{A} ^{B} ^{B} ^{C} ^{C}_{r}  _{b} _{r▷t} _{c}^{t} ^{r▷_{c}τ^{⋄}}^{ρ^{⋄}▷r}^{ev}^{ev} ^{r▷t}^{b▷(r▷t)}^{r▷t}

since the right branches under r(t ▷ t) [figure: ev] t in the outer diagrams have to agree.

The stacking conditions for R and T , and the independence of ρ_{⋄} and τ^{⋄} as well as ρ^{⋄} and τ_{⋄} imply that ρ^{⋄}s and rσ_{⋄} are modisms from R ▷ T to aR ▷ T . Again, naturality turns out to be a direct consequence of the modism axioms.

Just as in Section 2 coequalizers were necessary to define a meaningful composition of modules, we now expect equalizers to play the same role for right extensions of modules.

5.05. Assumption. Y locally has equalizers.

Again, for technical reasons we assume that a not necessarily canonical choice of equalizers has been made.

5.06. Proposition. Let b and c be endo-1-cells on B and C, respectively.

(0) The category ⟨b, c⟩mod has equalizers.

(1) If B and C are monads with carriers b and c respectively, then ⟨B, C⟩ (Y -mnd ) is closed under equalizers in ⟨b, c⟩mod .

Proof. (0) For modules S = ⟨σ_{⋄}, s, σ^{⋄}⟩ and V = ⟨ν_{⋄}, v, ν^{⋄}⟩ from b to c consider modisms ψ and χ from S to V . Let ⟨p, ξ⟩ be the equalizer of ψ and χ in [B, C] . This induces actions bp [figure: π_{⋄}] p and pc [figure: π^{⋄}] p via

[figure: pc; ξc; π^{⋄}; sc; ψc; χc; σ^{⋄}; vc; ν^{⋄}; p; ξ; s; ψ; χ; v; bp; bξ; π_{⋄}; bs; bψ; bχ; σ_{⋄}; bv; ν_{⋄}]

Since ξ is mono, we can derive the stacking condition for π_{⋄} and π^{⋄} directly from the stacking condition for σ_{⋄} and σ^{⋄}. Hence P = ⟨π_{⋄}, p, π^{⋄}⟩ together with ξ is the equalizer of ψ and χ in ⟨b, c⟩mod .

(1) The fact that ξ is mono implies ⟨π_{⋄}, p, π^{⋄}⟩ is compatible with the monad structures of A and B as specified in Proposition 3.04.(b).

5.07. Definition. For modules R = ⟨ρ_{⋄}, r, ρ^{⋄}⟩ from a to b and T = ⟨τ_{⋄}, t, τ^{⋄}⟩ from a to c define ⟨r ⋗ t, ξ⟩ to be the equalizer of ρ_{⋄} ▷ t and r ▷ τ_{⋄}^{▷} in ⟨b, c⟩mod . We denote the carrier of the module R ⋗ T by r ⋗ t, and the left and right actions by ρ^{⋄} ⋗ t and r ⋗ τ^{⋄}, respectively.

5.08. Theorem. All right extensions exist for modules between endo-1-cells in Y , and they are constructed by means of ⋗.

Proof. Consider modules a [figure: R] b and a [figure: T] c. Recall that R • (R ⋗ T) is defined in terms of the coequalizer r(r ⋗ t) [figure: ϑ] r • (r ⋗ t) of the 2-cells ρ^{⋄}(r ⋗ t) and r(ρ^{⋄} ⋗ t) from rb(r ⋗ t) to r(r ⋗ t) . By definition of ρ^{⋄} ▷ t and ρ^{⋄} ⋗ t we now have

A ^{B} ^{B} ^{C}_{r}  _{b} _{r⋗t}^{t}^{ρ^{⋄}}  ^{ξ}^{ev}^{r} ^{r▷t}  = ^{A} ^{B} ^{B} ^{C}_{r}  _{b} _{r⋗t}^{t}  ^{ξ}^{ρ^{⋄}▷t}^{ev} ^{r▷t}^{r▷t}  = ^{A} ^{B} ^{B} ^{C}_{r}  _{b} _{r⋗t}^{t}^{ρ^{⋄}⋗t}^{ξ}^{ev}^{r⋗t}^{r▷t}

Hence the transpose of r ⋗ t [figure: ξ] r ▷ t, i.e., the restriction of ev by ξ, coequalizes ρ^{⋄}(r ⋗ t) and r(ρ^{⋄} ⋗ t) . Consequently, there exists a unique 2-cell r • (r ⋗ t) [figure: EV] t that satisfies

A ^{B} ^{C}_{r} _{r⋗t}^{t} ^{ξ}^{ev}^{r▷t} = ^{A} ^{B} ^{C}_{r} _{r⋗t}^{t}^{ϑ}^{EV}^{r • (r⋗t)}  (5-02)

We wish to show that EV is a modism from R • (R ⋗ T) to T . Since aϑ is epi, EV commutes with the left actions:

A ^{A} ^{B} ^{C}_{a} _{r} _{r⋗t}^{t}^{τ_{⋄}} ^{ϑ}^{EV}^{r • (r⋗t)}^{t}  = ^{A} ^{A} ^{B} ^{C}_{a} _{r} _{r⋗t}^{t}  ^{ξ}^{ev}^{τ_{⋄}} ^{r▷t}^{t}  = ^{A} ^{A} ^{B} ^{C}_{a} _{r} _{r⋗t}^{t}  ^{ξ}^{ev}^{τ_{⋄}^{▷}}^{ev} ^{r▷t}^{t}^{a▷t}  = ^{A} ^{A} ^{B} ^{C}_{a} _{r} _{r⋗t}^{t}  ^{ξ}^{r▷τ_{⋄}^{▷}}^{ev}^{ev} ^{r▷t}^{r▷(a▷t)}^{a▷t}  =

= ^{A} ^{A} ^{B} ^{C}_{a} _{r} _{r⋗t}^{t}  ^{ξ}^{ρ_{⋄}▷t}^{ev}^{ev} ^{r▷t}^{(ar)▷t}^{a▷t}  = ^{A} ^{A} ^{B} ^{C}_{a} _{r} _{r⋗t}^{t}  ^{ξ}^{ρ_{⋄}} ^{ev}^{r} ^{r▷t}  = ^{A} ^{A} ^{B} ^{C}_{a} _{r} _{r⋗t}^{t}^{ρ_{⋄}} ^{ϑ}^{EV}^{r}^{r • (r⋗t)}  = ^{A} ^{A} ^{B} ^{C}_{a} _{r} _{r⋗t}^{t} ^{ϑ}^{ρ_{⋄} • (r⋗t)}^{EV}^{r • (r⋗t)}^{r • (r⋗t)}

And since ϑc is epi, EV commutes with the right actions as well:

A ^{B} ^{C} ^{C}_{r} _{r⋗t} _{c}^{t}^{ϑ}^{EV}^{τ^{⋄}}^{r • (r⋗t)}^{t}  = ^{A} ^{B} ^{C} ^{C}_{r} _{r⋗t} _{c}^{t}^{ξ}^{ev} ^{τ^{⋄}}^{r▷t}^{t}  = ^{A} ^{B} ^{C} ^{C}_{r} _{r⋗t} _{c}^{t}^{ξ}^{r▷_{v}τ^{⋄}}^{ev}^{r▷t}^{r▷t}  =

= ^{A} ^{B} ^{C} ^{C}_{r} _{r⋗t} _{c}^{t}^{r⋗τ^{⋄}}^{ξ}^{ev}^{r⋗t}^{r▷t}  = ^{A} ^{B} ^{C} ^{C}_{r} _{r⋗t} _{c}^{t}^{r⋗τ^{⋄}}^{ϑ}^{EV}^{r⋗t}^{r • (r⋗t)}  = ^{A} ^{B} ^{C} ^{C}_{r} _{r⋗t} _{c}^{t}^{ϑ}^{r • (r⋗τ^{⋄})}^{EV}^{r • (r⋗t)}^{r • (r⋗t)}

Pasting of modisms at R ⋗ T of EV corresponds to pasting of 2-cells at r ⋗ t of rξ ; ev in Y , cf. Diagram (5-02), and hence is injective. To establish the surjectivity, consider a modism R • S [figure: γ] T . Recall that R • S is defined in terms of the coequalizer φ of the modisms ρ^{⋄}s and rσ_{⋄} from RbS to RS. Let s [figure: ψ] r ▷ t be the transpose of φ ; γ. It suffices to show that ψ equalizes ρ_{⋄} ▷ t and r ▷ τ_{⋄}^{▷} in [B, C] . This follows from

A ^{A} ^{B} ^{C}_{a} _{r}  _{s}^{t} ^{ψ}^{ρ_{⋄}▷t}^{ev} ^{r▷t}^{(ar)▷t} = ^{A} ^{A} ^{B} ^{C}_{a} _{r}  _{s}^{t}^{ρ_{⋄}}  ^{ψ}^{ev}^{r} ^{r▷t}  = ^{A} ^{A} ^{B} ^{C}_{a} _{r}  _{s}^{t}^{ρ_{⋄}} ^{φ}^{γ}^{r}^{r • s}  = ^{A} ^{A} ^{B} ^{C}_{a} _{r}  _{s}^{t} ^{φ}^{ρ_{⋄} • s}^{γ}^{r • s}^{r • s}  =

= ^{A} ^{A} ^{B} ^{C}_{a} _{r}  _{s}^{t} ^{φ}^{γ}^{τ_{⋄}}^{r • s}^{r • s}  = ^{A} ^{A} ^{B} ^{C}_{a} _{r}  _{s}^{t} ^{ψ}^{ev}^{τ_{⋄}} ^{r▷t}^{t}  = ^{A} ^{A} ^{B} ^{C}_{a} _{r}  _{s}^{t} ^{ψ}^{r▷_{c}τ_{⋄}^{▷}}^{ev} ^{r▷t}^{(ar)▷t}

since the right branches of ar(ar ▷ t) [figure: ev] t in the outer diagrams agree.

5.09. Remark. If one drops the requirement of a tensor unit from the definition of a (not necessarily symmetric) left- (and right-) closed monoidal category, one arrives at the notion of a Lambek calculus. What we have here is a 2-dimensional analogue: if the base Y has all right extensions (and right liftings), the endo-1-cells, modules, and modisms nearly form a bicategory with the same property, except for the missing identity modules.

5.10. Theorem. The bicategories Y -mnd and Y -int have all right extensions. In Y -mnd they are formed via ⋗, just as for general modules, while in Y -int pre- and post-composition with the appropriate identitiy i-modules is additionally required.

Proof. For Y -mnd this is an immediate consequence of Proposition 5.06.(1).

If, on the other hand, A [figure: R] B and T [figure: T] C are i-modules between interpolads, the actions ρ^{⋄} ⋗t and r ⋗τ^{⋄} of R⋗T may fail to have the coequalizer property required in axioms (I0) and (I1), although the corresponding diagrams (cf. (3-00)) are guaranteed to commute, since ξ is mono.

Let b(r ⋗ t) [figure: φ] b • (r ⋗ t) be the chosen coequalizer of β(t ⋗ t) and b(ρ^{⋄} ⋗ t) . There exists a unique modism b • (r ⋗ t) [figure: μ] r ⋗ t that satisfies φ ; μ = ρ^{⋄} ⋗t. Similarly, for the coequalizer (r ⋗ t)c [figure: ψ] (r ⋗ t) • c of (r⋗τ^{⋄})c and (r⋗t)ζ we find a uniqe modism (r ⋗ t) • c [figure: ν] r ⋗ t with ψ ; ν = r⋗τ^{⋄}. These clearly satisfy (b • ν) ; μ = (μ • c) ; ν. Let η be the resulting modism from the i-module B • (R ⋗ T) • C to R ⋗ T , cf., Proposition 3.03.. We see that B • (R ⋗ T) • C together with (r • η) ; EV constitutes the desired right extension in Y -int .

5.11. Examples. (0) Consider modules R = ⟨ρ_{⋄}, R, ρ^{⋄}⟩ from A to B and T = ⟨τ_{⋄}, T, τ^{⋄}⟩ from A to C in the suspension of set (cf. Example 3.07.(0)). The set R ⋗ T ⊆ R ▷ T = T^{R} consists of left-homomorphisms, i.e., functions R [figure: f] T that satisfy ρ_{⋄} ; f = (A × f) ; τ_{⋄}. The left action ρ^{⋄} ⋗ T maps ⟨b, f⟩ ∈ R × (R ⋗ T) to the left-homomorphism R [figure: f_{b}] T defined by rf_{b} = (rb)f , and the right action R ⋗ τ^{⋄} maps ⟨f, c⟩ ∈ (R ⋗ T) × C to the left-homomorphism R [figure: f^{c}] T defined by rf^{c} = ((rf)) c. The restriction of the ordinary evaluation function R × (R ▷ T) [figure: ev] T to R × (R ⋗ T) coequalizes ρ^{⋄} × (R ⋗ t) and R × (ρ^{⋄} ⋗ T) , which induces the function R • (R ⋗ T) [figure: EV] T .

(1) Since for Y = rel local equalizers are trivial, the operators ▷ and ⋗ agree. Given modules a [figure: r] b and a [figure: t] c, their right extension is given by

r ▷ t = { ⟨y, z⟩ ∈ B × C | ∀x ∈ A. ⟨x, y⟩ ∈ r ⇒ ⟨x, z⟩ ∈ t }  (5-03)

and the 2-cell ev is the inclusion r(r ▷ t) ⊆ t. Observe that (rb) ⊆ r implies b(r ▷ t) ⊆ r ▷ t, and that tc ⊆ t implies (r ▷ t)c ⊆ r ▷ t. If r and t are i-modules between idempotent relations, r ⋗ t may fail to be idempotent. Then the relation product b(r ⋗ t)c together with the inclusion rb(r ▷ t)c = r(r ▷ t)c ⊆ r(r ▷ t) ⊆ t serves as the right extension, cf., Rosebrugh and Wood [8].

(2) As we indicated in Example 3.07.(3), the case of rel is subsumed by that of quantaloids, i.e., categories enriched in cslat . Given modules a [figure: r] b and a [figure: t] c, (the underlying 1-cell of) their right extension according to the first adjunction ind (5-00) is given by

r ▷ t = sup{ s ∈ [B, C] : rs ≤ t }  (5-04)

and this coincides with r ⋗ t. Equation (5-03) above expresses the same fact at the level of elements of the sets B and C. The description of the right extension of i-modules is as in part (1) above. If we view cslat as self-enriched, interpolads in this context are additive idempotent endo-functions, while monads correspond to additive closure operators.

(3) In the bicategory spn (cf. Example 3.07.(4)) the right extension along A [figure: r]B of A [figure: t]C consists of the matrix r ▷ t with ⟨y, z⟩-component ∏_{x∈A}⟨x, y⟩r ▷ ⟨x, z⟩t and the evaluation

∑_{y∈B}(⟨x, y⟩r × ∏_{u∈A} ⟨u, y⟩r ▷ ⟨u, z⟩t) [figure: ev] ⟨x, z⟩t

that feeds h ∈ ⟨x, y⟩r into the x-component of an A-tuple of functions. The ⟨y, z⟩-component of the right extension along a module A [figure: R]B of A [figure: T]C is the equalizer of the two canonical functions from ∏_{x∈A} ⟨x, y⟩r ▷ ⟨x, z⟩t to

∏_{x∈A}(∑_{u∈A} ⟨x, u⟩a × ⟨u, y⟩r) ▷ ⟨x, z⟩t ≅ ∏_{u∈A} ⟨u, y⟩r ▷ (∏_{x∈A} ⟨x, u⟩a ▷ ⟨x, z⟩t)

Concretely this means that the formal arrows B ∋ y [figure: s] z ∈ C are A-indexed families of functions ⟨x, y⟩r [figure: s_{x}] ⟨x, z⟩t that satisfy the following compatibility condition: if in the following diagram of formal arrows the left triangle commutes, so does the right one:

[figure: x; hs_{x}; h; f; y; z; u; g; gs_{u}]

This also describes right extensions in prof , while in tax = spn -int we still need to pre- and post-compose with the appropriate identity i-modules.

Acknowledgement. I am grateful to Jirí Adámek, Victor Pollara and Werner Struckmann at the TU Braunschweig for inspiring discussions, and to the anonymous referee for helpful comments. Also I would like to thank Kristoffer H. Rose and Ross Moore for their superb macro package XY-pic 3.2 that made the diagrams in this paper possible.

References

[1] Benabou, J. Introduction to bicategories. In Reports of the Midwest Category Seminar. I, no. 47 in Lecture Notes in Mathematics. Springer-Verlag, Berlin – New York, 1967, pp. 1–77.

[2] Borceux, F. Handbook of Categorical Algebra 1, Basic Category Theory, vol. 50 of Encyclopedia of Mathematics and its Applications. Cambridge University Press, Cambridge, 1994.

[3] Carboni, A., Kasangian, S., and Walters, R. An axiomatics for bicategories of modules. J. Pure Appl. Algebra 45 (1987), 127–141.

[4] Gordon, R., Power, A. J., and Street, R. Coherence for Tricategories. No. 558 in Memoirs of the AMS. AMS, Providence, RI, 1995.

[5] Hayashi, S. Adjunction of semifunctors: Categorical structures in non-extensional lambda-calculus. Theor. Comp. Sci. 41 (1985), 95–104.

[6] Hoofman, R., and Moerdijk, I. A remark on the theory of semi-functors. Math Struct. in Comp. Science 5 (1995), 1–8.

[7] Lawvere, F. W. Metric spaces, generalized logic, and closed categories. Rend. Sem. Mat. Fis. Milano 43 (1973), 135–166.

[8] Rosebrugh, R., and Wood, R. J. Constructive complete distributivity IV. Appl. Cat. Struct. 2 (1994), 119–144.

[9] Street, R. Enriched categories and cohomology. Quaestiones Math. 6 (1983), 265– 283.

[10] Street, R., and Joyal, A. The geometry of tensor calculus I. Advances in Mathematics 88 (1991), 55–112.

[11] Street, R., and Walters, R. F. C. Yoneda structures on 2-categories. Journal of Algebra 50 (1978), 350–379.

[12] Vickers, S. Information systems for continuous posets. Theoretical Computer Science 114 (1993), 201–229.

Institut für Theoretische Informatik TU Braunschweig, P.O. Box 3329 38023 Braunschweig, Germany

Email: koslowjiti.cs.tu-bs.de

This article may be accessed via WWW at http://www.tac.mta.ca/tac/ or by anony-

mous ftp at ftp.tac.mta.ca/pub/tac/html/volumes/1997/n8/n8.{dvi,ps}

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods.

Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication.

The method of distribution of the journal is via the Internet tools WWW/ftp. The journal is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive (by e-mail) abstracts of articles as they are published. Full text of published articles is available in .dvi and Postscript format. Details will be e-mailed to new subscribers and are available by WWW/ftp. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX is the preferred flavour. TEX source of articles for publication should be submitted by e-mail directly to an appropriate Editor. They are listed below. Please obtain detailed information on submission format and style files from the journal’s WWW server at URL http://www.tac.mta.ca/tac/ or by anonymous ftp from ftp.tac.mta.ca in the directory pub/tac/info. You may also write to tac@mta.ca to receive details by e-mail.

Editorial board.

John Baez, University of California, Riverside: baez@math.ucr.edu Michael Barr, McGill University: barr@triples.math.mcgill.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: r.brown@bangor.ac.uk Jean-Luc Brylinski, Pennsylvania State University: jlb@math.psu.edu Aurelio Carboni, Università della Calabria: carboni@unical.it P. T. Johnstone, University of Cambridge: ptj@pmms.cam.ac.uk G. Max Kelly, University of Sydney: kelly m@maths.su.oz.au Anders Kock, University of Aarhus: kock@mi.aau.dk F. William Lawvere, State University of New York at Buffalo: wlawverer@acsu.buffalo.edu Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.ruu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@cs.dal.ca Andrew Pitts, University of Cambridge: ap@cl.cam.ac.uk Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz James Stasheff, University of North Carolina: jds@charlie.math.unc.edu Ross Street, Macquarie University: street@macadam.mpce.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Sydney: walters b@maths.su.oz.au R. J. Wood, Dalhousie University: rjwood@cs.da.ca